For a monopolistic firm, its demand is P = 100 - Q while MR = 100 - 2Q, if its MC = 14, how much it should produce to maximize its profit?
Answer #1

P = 100 - Q

MR = 100 - 2Q

MC = 14

Profit maximizing condition

MR = MC

100 - 2Q = 14

100 - 14 = 2Q

86 = 2Q

Q = 86/2

= 43  

P = 100 - 43  

= 57